# Second- and third-order properties of multidimensional Langevin equations

### v9

@@ Abstract.
Recent work has addressed the problem of inferring Langevin dynamics from data. In this work, we address the problem of relating terms in the Langevin equation to statistical properties, such as moments of the probability density function and of the probability current density, as well as covariance functions. We first review the case of linear Gaussian dynamics, and then consider extensions beyond this simple case. We address the question of quantitative significance of effects. We also analyze underdamped (second-order) processes, specifically in the limit where dynamics in state space is almost Markovian.
@@ 1. Introduction
Finally, we consider underdamped (second-order) processes, specifically in the limit where dynamics in state space is almost Markovian.
@@ 2.1. Preliminaries: Definition, covariance functions, and detailed balance
with $\mathbf{A}$ and $\mathbf{D}$ constant, and $\delta(\cdot)$ is Dirac delta function. The covariance matrix $\mathbf{C}:=\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ is determined from the Lyapunov equation:
@@
where $\mathds{1}$ denotes the identity matrix of appropriate dimension. When comparing to experimental data, this relation is not to be considered as a test of the linear model, as it is a simple consequence of stationarity of $\mathbf{x}$ and the model fit, by Itô’s lemma:
@@
upon dividing by $\tau$ and taking $\tau\to 0$, where we used $\langle\mathbf{x}(\tau)\mathbf{x}(\tau)^{\mathsf{T}}\rangle=\langle\mathbf{x}(0)\mathbf{x}(0)^{\mathsf{T}}\rangle$. (In certain cases, care must be taken when evaluating these expressions; see Appendix B.)
Of interest is the covariance function:
@@
where $\tau\geq 0$ is always assumed. Note that this is satisfied even in the case of non-constant diffusion.
Of interest also is the condition for detailed balance. A requirement for detailed balance is that the covariance function is symmetric, $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle=\langle\mathbf{x}(0)\mathbf{x}(\tau)^{\mathsf{T}}\rangle$. By differentiating with respect to $\tau$, we get $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$. To gain insight into this relation, we switch to “covariance-identity” coordinates in which $\mathbf{C}=\mathds{1}$ [13]. In such a basis, $\mathbf{A}$ is symmetric, which means it is diagonalizable by an orthogonal basis. In these coordinates, an orthogonal basis means uncorrelated components. By the Lyapunov equation, $\mathbf{D}$ must also be diagonal in this basis. Hence the system decomposes into a number of noninteracting one-dimensional linear systems [4]. Ref. [4] also contains good discussions about general properties of linear Gaussian systems.
The Lyapunov equation implies a relationship between the eigenvalues of $\mathbf{A}$ and of $\mathbf{D}\mathbf{C}^{-1}$. We work in a coordinate system in which $\mathbf{C}=\mathds{1}$. Let $\lambda$ be an eigenvalue of $\mathbf{A}$ and $\mathbf{v}$ be a corresponding eigenvector. Then:
(7) $\mathbf{v}^{\mathsf{H}}\mathbf{D}\mathbf{v}=-\mathbf{v}^{\mathsf{H}}\left({\frac{\mathbf{A}+\mathbf{A}^{\mathsf{T}}}{2}}\right)\mathbf{v}=-(\Re\lambda)(\mathbf{v}^{\mathsf{H}}\mathbf{v}),$
therefore $-\Re\lambda$ is bounded by the smallest and largest eigenvalues of $\mathbf{D}$ [14].
@@ 2.2. Comparison of experimental data with theoretical predictions
Now, we turn to the question of comparing experimental data and theoretical predictions under a linear model. We may consider comparing the deviation of $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ to $\mathbf{C}$. That is, if we consider a whitened dataset for which $\mathbf{C}=\mathds{1}$, the entries of the deviation matrix in these coordinates should be compared to unity. For a coordinate-independent description, we may look at the symmetric and antisymmetric parts of the deviation matrix and multiply these on the right by $\mathbf{C}^{-1}$, and inspect the eigenvalues, which are purely real and imaginary in the symmetric and antisymmetric cases, respectively.
We denote the deviation $\mathbf{M}:=\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle_{\text{expt}}-\mathbf{C}_{\text{theo}}$. From now on we suppress the “theo” subscript and treat $\mathbf{C}$ as fixed. We may consider the eigenvalues of $\mathbf{M}\mathbf{C}^{-1}$, which are real since we have a matrix that is symmetric in coordinates in which $\mathbf{C}=\mathds{1}$. We may imagine an ensemble of stochastic systems, in which deviations of $\mathbf{M}$ from $\mathbf{0}$ are permitted. We may ask the question of what a suitable variance of $\mathbf{M}$ would be. We may then compare the elements of $\mathbf{M}$ to the square root of the “ensemble variance”. For biological applications, a factor of 0.3 or more may be considered significant.
@@
where $\sim$ denotes ensemble expectation. We may treat $(i,j)$ as a single index and similarly for $(i^{\prime},j^{\prime})$, and multiply the l.h.s. (left-hand side) by the matrix inverse of the r.h.s. (right-hand side). We denote by $\boldsymbol{\Gamma}$ the resulting matrix:
@@
It so happens that $\boldsymbol{\Gamma}$ has at most one non-zero eigenvalue $\mu$, which is equal to its trace. (In future sections, this will not necessarily be the case.) We have:
@@
Now if we neglect permutations in $(i,j)$ and $(i^{\prime},j^{\prime})$, we can see that the matrix product of $M^{ij}M^{i^{\prime}j^{\prime}}$ with the matrix inverse of the quantity following $v_{ij}$ on the r.h.s. of Eq. (15) has the same eigenvalue $\mu$. Thus we may consider instead of Eq. (8) the relation:
@@
Now we turn to covariance functions with a time-lag. We may again denote the deviation $\mathbf{M}:=\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle_{\text{expt}}-\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle_{\text{theo}}$. We can separate $\mathbf{M}$ into symmetric and antisymmetric components:
@@
(24) $T^{ij}T^{i^{\prime}j^{\prime}}$ $\sim\frac{C^{ii^{\prime}}C^{jj^{\prime}}-C^{ij^{\prime}}C^{ji^{\prime}}}{2}.$
@@ 2.3. Probability current density and angular momentum
The stochastic rotation frequency is bounded above by the geometric mean of the relaxation rates, $\left|\omega_{\mathrm{stoch}}\right|\leq\sqrt{\lambda_{x}\lambda_{y}}$, with equality attained when $\mathbf{D}$ is singular ($D^{xx}D^{yy}=(D^{xy})^{2}$). Now, we turn to the case where $\mathbf{A}$ has a pair of complex conjugate eigenvalues. Without loss of generality, after a change of coordinates, we may write:
@@
It can be seen from elementary algebra that $\omega_{\mathrm{stoch}}\leq\sqrt{\lambda^{2}+\omega^{2}}$, with equality holding if and only if $D^{xx}D^{yy}=0$, i.e., the matrix $\mathbf{D}$ is singular.
@@
Because the phase-space velocity $\mathbf{J}(\mathbf{x})/p(\mathbf{x})$ is linear, $\langle{}\circ\dot{\phi}\mid r,\phi\rangle$ is independent of $r$. Also, by our choice of coordinates, the distribution of $\mathbf{x}$ is isotropic, which means that $r$ is independent of $\phi$. Thus $L=\langle r^{2}\circ\dot{\phi}\rangle=\langle r^{2}\rangle\langle\dot{\phi}\rangle=2\langle\dot{\phi}\rangle$, as desired. We have numerically verified this for a system obeying Eq. (32) where $\lambda=\omega=1$, $D^{xx}=1$, $D^{yy}=10$. We simulated this system using Euler–Maruyama discretization with $\Delta t=0.005$ for 1000 time-units. We see that the measured stochastic rotation frequency is in agreement with the theoretical value of 1.23 (Fig. 1).
@@
If $h$ is at least comparable to unity, then we consider that broken detailed balance is significant. (We may understand the factor of 2 in the definition of $\mathbf{H}$ by considering that the symmetric part of $\mathbf{A}\mathbf{C}$ is $-\mathbf{D}$ by the Lyapunov equation, so the antisymmetric part should be compared to $\mathbf{D}$.) We may illustrate this on the two examples above. For two real eigenvalues of $\mathbf{A}$,
@@
For the case of complex conjugate eigenvalues of $\mathbf{A}$,
@@
For angular momenta of $d$-dimensional systems, we may follow the part of the previous discussion pertaining to the antisymmetric part, except with $\mathbf{C}$ replaced by $2\mathbf{D}$, i.e.:

### v11

@@ Abstract.
Recent work has addressed the problem of inferring Langevin dynamics from data. In this work, we address the problem of relating terms in the Langevin equation to statistical properties, such as moments of the probability density function and of the probability current density, as well as covariance functions. We first review the case of linear Gaussian dynamics, and then consider extensions beyond this simple case. We address the question of quantitative significance of effects. We also analyze underdamped (second-order) processes, specifically in the limit where dynamics in state space is almost Markovian. Finally, we address detection of non-Markovianity.
@@ 1. Introduction
Finally, we consider underdamped (second-order) processes, specifically in the limit where dynamics in state space is almost Markovian, as well as detection of non-Markovianity.
@@ 2.1. Preliminaries: Definition, covariance functions, and detailed balance
 [^1] We have omitted a possible additive constant in $\dot{\mathbf{x}}$. In the stationary case, a system with an additive constant may be transformed into Eq. (1) by considering instead the variable $\mathbf{x}-\mathbf{c}$, where $\mathbf{c}$ is a constant (in this case, $\mathbf{c}=\langle\mathbf{x}\rangle$). This simplification is used throughout.
where $\boldsymbol{\xi}(t)$ is a Gaussian random term, $\mathbf{A}$ and $\mathbf{D}$ are constant, and $\delta(\cdot)$ is Dirac delta function. The covariance matrix $\mathbf{C}:=\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ is determined from the Lyapunov equation [13]:
@@
where $\mathds{1}$ denotes the identity matrix of appropriate dimension, and $\otimes$ denotes tensor or Kronecker product [^2] For a matrix representing a linear transformation (i.e., a second-rank mixed tensor), the first index is a superscript (contravariant) and the second index is a subscript (covariant). Thus the $(i,j)$ element of such a matrix $A$ is denoted $A^{i}_{j}$. The tensor or Kronecker product $A\otimes B$ of two second-rank mixed tensors $A$ and $B$ is defined as $(A\otimes B)^{ij}_{kl}=A^{i}_{k}B^{j}_{l}$. Tensors with equal number of contravariant and covariant indices may be regarded as matrices where all the contravariant (respectively covariant) indices are regarded as a single “super-index”. The usual matrix operations can then be defined in this way, e.g. multiplication as $(AB)^{ij}_{kl}:=A^{ij}_{i^{\prime}j^{\prime}}B^{i^{\prime}j^{\prime}}_{kl}$ and inversion correspondingly, resulting in tensors of the same type. For inversion, we may argue as follows. Suppose $A$ and $B$ are tensors of the same type with equal number of contravariant and covariant indices, and suppose they are matrix inverses of each other in a particular coordinate system. Because the equalities $AB=\mathds{1}$ and $BA=\mathds{1}$ are invariant under a change of basis, we conclude that $A$ and $B$ are matrix inverses of each other in any coordinate system, hence $A^{-1}=B$ transforms according to the same law as $A$. On the other hand, the covariance and diffusion matrices are second-rank contravariant tensors and thus their matrix inverses represent covariant tensors. We can explicitly compute the transformation law for such a matrix under a change of basis matrix $\mathbf{R}$. A second-rank contravariant tensor $\mathbf{C}$ transforms into $\widetilde{\mathbf{C}}$ where $\widetilde{C}^{ij}=(R^{-1})^{i}_{k}(R^{-1})^{j}_{l}C^{kl}=(\mathbf{R}^{-1}\mathbf{C}(\mathbf{R}^{-1})^{\mathsf{T}})^{ij}$. Its matrix inverse $\mathbf{C}^{-1}$ thus transforms into $\widetilde{C}^{-1}_{ij}=(\mathbf{R}^{\mathsf{T}}\mathbf{C}^{-1}\mathbf{R})_{ij}=R^{k}_{i}R^{l}_{j}C^{-1}_{kl}$, which is the transformation law for a covariant tensor. We may establish similar rules for tensors with an even number of indices, regardless of type. Also, symmetry or antisymmetry of tensors under exchange of indices that are either both contravariant or both covariant is invariant under a change of basis.. When comparing to experimental data, this relation is not to be considered as a test of the linear model, as it is a simple consequence of stationarity of $\mathbf{x}$ and the model fit, by Itô’s lemma:
@@
upon dividing by $\tau$ and taking $\tau\to 0$, where we used $\langle\mathbf{x}(\tau)\mathbf{x}(\tau)^{\mathsf{T}}\rangle=\langle\mathbf{x}(0)\mathbf{x}(0)^{\mathsf{T}}\rangle$. Eq. (4) follows if the derivative can be taken inside the expectation (however, see Appendix B for a counterexample). This type of manipulation will become useful later.
Of interest is the covariance function [3]:
@@
where $\tau\geq 0$ is always assumed. This can be derived by solving the differential equation for the conditional expectation:
(7) $\frac{\mathrm{d}}{\mathrm{d}\tau}\langle\mathbf{x}(\tau)\mid\mathbf{x}(0)\rangle=\mathbf{A}\langle\mathbf{x}(\tau)\mid\mathbf{x}(0)\rangle,$
 [^3] For the stationary case, we require that the probability distribution approaches the stationary distribution as time passes. This means that the eigenvalues of $\mathbf{A}$ must have strictly negative real parts. Additionally, non-vanishing real parts of eigenvalues of $\mathbf{A}$ are required for $\mathbf{A}\otimes\mathds{1}+\mathds{1}\otimes\mathbf{A}$ to be invertible. We may demonstrate this as follows. Let $\mathbf{v}$ be an eigenvector of $\mathbf{A}$ with eigenvalue $\lambda+i\omega$ ($\lambda,\omega$ real). Then its complex conjugate $\mathbf{v}^{*}$ is an eigenvector of $\mathbf{A}$ with eigenvalue $\lambda-i\omega$. We see that $\mathbf{v}\otimes\mathbf{v}^{*}$ is an eigenvector of $\mathbf{A}\otimes\mathds{1}+\mathds{1}\otimes\mathbf{A}$ with eigenvalue $2\lambda$. Thus if $\mathbf{A}\otimes\mathds{1}+\mathds{1}\otimes\mathbf{A}$ is invertible, we must have $\lambda\neq 0$.
and multiplying on the right by $\mathbf{x}(0)^{\mathsf{T}}$ and using the law of iterated expectations, again assuming that the derivative can be taken inside the expectation. Note that this result does not depend on homogeneity of diffusion.
Of interest also is the condition for detailed balance. A requirement for detailed balance is that the covariance function is symmetric, $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle=\langle\mathbf{x}(0)\mathbf{x}(\tau)^{\mathsf{T}}\rangle$. By differentiating with respect to $\tau$, we get $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$ [^4] By expanding $e^{\mathbf{A}\tau}$ in a Taylor series, it is seen that this condition implies symmetry of the covariance function.. To gain insight into this relation, we switch to “covariance-identity” coordinates in which $\mathbf{C}=\mathds{1}$ [13]. (We assume that $\mathbf{C}$ is finite and positive definite, so that such a transformation can be made.) In these coordinates, detailed balance implies that $\mathbf{A}$ is symmetric, which means it is diagonalizable by an orthogonal basis. In “covariance-identity” coordinates, an orthogonal basis means uncorrelated components (because $\langle(\mathbf{v}^{\mathsf{T}}\mathbf{x})(\mathbf{x}^{\mathsf{T}}\mathbf{w})\rangle=\mathbf{v}^{\mathsf{T}}\mathbf{w}$ for constant vectors $\mathbf{v}$, $\mathbf{w}$). By the Lyapunov equation, $\mathbf{D}$ must also be diagonal in this basis. Hence the system decomposes into a number of noninteracting one-dimensional linear systems [4]. Ref. [4] also contains good discussions about general properties of linear Gaussian systems.
The Lyapunov equation implies a relationship between the eigenvalues of $\mathbf{A}$ and of $\mathbf{D}\mathbf{C}^{-1}$. (These are second-rank mixed tensors, whose eigenvalues are independent of the basis in which they are represented.) We again work in a coordinate system in which $\mathbf{C}=\mathds{1}$. Let $\lambda$ be an eigenvalue of $\mathbf{A}$ and $\mathbf{v}$ be a corresponding eigenvector. Then:
(8) $\mathbf{v}^{\mathsf{H}}\mathbf{D}\mathbf{v}=-\mathbf{v}^{\mathsf{H}}\left({\frac{\mathbf{A}+\mathbf{A}^{\mathsf{T}}}{2}}\right)\mathbf{v}=-(\Re\lambda)(\mathbf{v}^{\mathsf{H}}\mathbf{v})$
($\mathsf{H}$ denoting Hermitian conjugate and $\Re$ denoting the real part), therefore $-\Re\lambda$ is bounded by the smallest and largest eigenvalues of $\mathbf{D}$ [14].
@@ 2.2. Comparison of experimental data with theoretical predictions
Now, we turn to the question of comparing experimental data and theoretical predictions under a linear model. As stated in the introduction, we may not necessarily care about statistically resolvable deviations if they are quantitatively small. We thus desire a way to evaluate the size of deviations. We may consider comparing the deviation between experiment and theory of $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ to $\mathbf{C}$. That is, if we consider a whitened dataset for which $\mathbf{C}=\mathds{1}$, the entries of the “deviation matrix” (i.e., experiment minus theory of $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$) in these coordinates should be compared to unity. For a coordinate-independent description, we may look at the symmetric and antisymmetric parts of the deviation matrix and multiply these on the right by $\mathbf{C}^{-1}$, and inspect the eigenvalues, which are purely real and imaginary in the symmetric and antisymmetric cases, respectively.
We denote the deviation $\mathbf{M}:=\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle_{\text{expt}}-\mathbf{C}_{\text{theo}}$, where “expt” and “theo” subscripts denote experimental and theoretical values, respectively. From now on we suppress the “theo” subscript and treat $\mathbf{C}$ as fixed. We may consider the eigenvalues of $\mathbf{M}\mathbf{C}^{-1}$, which are real because we have a matrix (again, a second-rank mixed tensor) that is symmetric in coordinates in which $\mathbf{C}=\mathds{1}$. We may imagine an ensemble of stochastic systems, each having its own value of $\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ which may differ from $\mathbf{C}$, and again denote the difference by $\mathbf{M}$. We assume that the ensemble mean of $\mathbf{M}$ is zero, and we now ask the question of what a suitable ensemble variance of $\mathbf{M}$ would be that reflects significant deviations. We may then compare the elements of the experimental value of $\mathbf{M}$ to the square root of the ensemble variance. For biological applications, a factor of 0.3 or more may be considered significant.
@@
where $\sim$ denotes ensemble expectation. We may treat $(i,j)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime})$, and multiply the l.h.s. (left-hand side) by the matrix inverse of the r.h.s. (right-hand side). We denote by $\boldsymbol{\Gamma}$ the resulting matrix:
@@
It so happens that $\boldsymbol{\Gamma}$ has at most one non-zero eigenvalue $\mu$, which is equal to its trace. (This is because in coordinates where $\mathbf{C}=\mathds{1}$, the matrix $\boldsymbol{\Gamma}$ takes the form $\mathbf{v}\mathbf{v}^{\mathsf{T}}$, and therefore has rank at most 1.) We have:
@@
Now if we neglect permutations in $(i,j)$ and $(i^{\prime},j^{\prime})$ [^5] I.e., only one of $(i,j)$ and $(j,i)$ is included in the list of “super-indices” (and similarly for $(i^{\prime},j^{\prime})$). This is necessary to invert $(C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}})/2$, since this quantity is invariant upon swapping $i\leftrightarrow j$ (or $i^{\prime}\leftrightarrow j^{\prime}$)., we can see that the matrix product of $M^{ij}M^{i^{\prime}j^{\prime}}$ with the matrix inverse of $(C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}})/2$ has left eigenvector $v_{ij}$ and eigenvalue $\mu$. Thus we may consider instead of Eq. (9) the relation:
@@
Now we turn to covariance functions with a time-lag. We may again denote the deviation $\mathbf{M}:=\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle_{\text{expt}}-\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle_{\text{theo}}$. We can separate $\mathbf{M}$ into symmetric and antisymmetric components, denoted by $\mathbf{S}$ and $\mathbf{T}$ respectively:
@@
(25) $T^{ij}T^{i^{\prime}j^{\prime}}$ $\sim\frac{C^{ii^{\prime}}C^{jj^{\prime}}-C^{ij^{\prime}}C^{ji^{\prime}}}{2},$
(26) $S^{ij}T^{i^{\prime}j^{\prime}}$ $\sim 0.$
@@ 2.3. Probability current density and angular momentum
The stochastic rotation frequency is bounded above by the geometric mean of the relaxation rates, $\left|\omega_{\mathrm{stoch}}\right|\leq\sqrt{\lambda_{x}\lambda_{y}}$, with equality attained if and only if $\mathbf{D}$ is singular ($D^{xx}D^{yy}=(D^{xy})^{2}$). Now, we turn to the case where $\mathbf{A}$ has a pair of complex conjugate eigenvalues. Without loss of generality, after a change of coordinates, we may write:
@@
It can be seen from elementary algebra that $\omega_{\mathrm{stoch}}\leq\sqrt{\lambda^{2}+\omega^{2}}$, with equality holding if and only if $\mathbf{D}$ is singular ($D^{xx}D^{yy}=0$).
@@
Because the phase-space velocity $\mathbf{J}(\mathbf{x})/p(\mathbf{x})$ is linear, $\langle{}\circ\dot{\phi}\mid r,\phi\rangle$ is independent of $r$ (see section 4 for a derivation). Also, by our choice of coordinates, the distribution of $\mathbf{x}$ is isotropic, which means that $r$ is independent of $\phi$. Thus $L=\langle r^{2}\circ\dot{\phi}\rangle=\langle r^{2}\rangle\langle\dot{\phi}\rangle=2\langle\dot{\phi}\rangle$, as desired. We have numerically verified this for a system obeying Eq. (34) where $\lambda=\omega=1$, $D^{xx}=1$, $D^{yy}=10$. We simulated this system using Euler–Maruyama discretization with $\Delta t=0.005$ for 1000 time-units. We see that the measured stochastic rotation frequency is in agreement with the theoretical value of 1.23 (Fig. 1).
@@
If $h$ is at least comparable to unity, then we consider that broken detailed balance is significant. (We may understand the factor of 2 in the definition of $\mathbf{H}$ by considering that the symmetric part of $\mathbf{A}\mathbf{C}$ is $-\mathbf{D}$ by the Lyapunov equation, so the antisymmetric part should be compared to $\mathbf{D}$.) We may illustrate this on the two examples above. For two real eigenvalues $-\lambda_{x}$ and $-\lambda_{y}$ of $\mathbf{A}$,
@@
For the case of complex conjugate eigenvalues $-\lambda\pm i\omega$ of $\mathbf{A}$ with $D^{xy}=0$,
@@
For angular momenta of $d$-dimensional systems, we may follow the part of the previous discussion pertaining to the antisymmetric part, except with $\mathbf{C}$ replaced by $2\mathbf{D}$, i.e., for the “ensemble covariance”: